Let u=<-1,3> and v=<4,-6>.
Find each of the following:
a. 3u-2v
b. The normalization of v

Answer by stanbon(48564)

Let u=<-1,3> and v=<4,-6>.
Find each of the following:
a. 3u-2v = 3(-1,3)-2(4,-6)=(-3-8.9+12)=(-11,21)
b. The normalization of v:
|v|= sqrt(4^2+6^2) = sqrt52 = 2sqrt13
So, norm of v = (4/(2sqrt13),-6/(2sqrt13))
===============
Cheers,
Stan H.